# Large Differences Between Stochastic and Deterministic Kinetics in a Simple Autocatalytic Reaction Network

### v3

@@ I Introduction
For all of the above phenomena, the first-order process (1) is indispensable, and therefore autocatalysis (2) alone is not sufficient as a minimal effective model. For example, although a simple autocatalytic reaction (2) can be used to model the aggregation of a misfolded protein molecules Arslan and Laurenzi (2008), the reaction (1) is necessary to describe the misfolding process.
The deterministic rate equations for reactions (1)–(2) are straightforward to solve and are used to fit experimental data on both nanoparticle formation Watzky and Finke (1997) and protein aggregation in vitro Watzky et al. (2008); Morris et al. (2008, 2009); Iashchishyn et al. (2017). However, some applications of the FWM may involve relatively small systems—e.g., modeling protein misfolding in vivo or disease spread in small populations—where stochastic approaches such as CME or SSA are more appropriate. The existence of significant differences between stochastic and deterministic descriptions of this model further motivates deriving an analytical solution of the CME.
@@ II.1 Irreversible reactions
The general analytical solution (10)–(11) of Eqs. (6) has appeared in various contexts and has been derived independently by different authors Allnatt and Jacobs (1968); Arslan and Laurenzi (2008); Lee and Kim (2012); Rudnicki (2014); Jedrak (2014). In contrast, by "explicit analytical solution" we mean here not the general form (10)–(11) of the solution, but the most concise analytical form of both $P_{n}(t)$ and its lowest moments that can be obtained for a specific choice of the coefficients $r_{n}$.
@@ III.1.1 Probability distribution and its two lowest moments
Analytical expressions for the moments (III.1.1)–(III.1.1) reduce the double sum (which arises when calculating the moment directly from its definition using $P_{n}(t)$ given by Eq. 22) to a single sum, lowering the computational cost. A caveat is that numerical evaluation of Eqs. (III.1.1) and (III.1.1) requires arbitrary-precision arithmetic: for $M=100$ the absolute values of the coefficients $E^{(1)}_{l}$ and $E^{(2)}_{l}$ reach $10^{30}$ and $10^{33}$, respectively, so small relative round-off errors at short times can cause large absolute errors in the moments. [^3] Here, the GNU Multiple Precision Floating-Point Reliable Library (GNU MPFR) was used for this purpose (https://www.mpfr.org). In fact, instead of relying on the potentially cumbersome arbitrary-precision evaluation of Eqs. (III.1.1)–(III.1.1), one can simply solve the set of $N+1$ ODEs in Eq. (6) by direct numerical integration. We performed the calculations using both approaches and obtained perfect agreement with the analytical solution.
@@ III.1.3 Qualitative explanation of the observed effects
To study the stochastic kinetics of the Finke–Watzky model, one may replace the full CME analysis with an approximate approach based on a truncated system of moment equations. Here we examine such equations in order to provide a partial explanation of the stochastic delay phenomenon. At the same time, our results highlight the limitations of moment-equation approaches when it comes to accurate quantitative prediction of the magnitude of differences between stochastic and deterministic kinetics.
@@
where $\tilde{M}=M+k_{1}/k_{2}$. Note that we have
@@
The former inequality is immediate (as $0\leq a\leq\tilde{M}$), while the latter follows from the relations $\tilde{M}\,\mu_{1A}(t)\geq M\,\mu_{1A}(t)\geq N\,\mu_{1A}(t)$ and $N\,\mu_{1A}(t)\geq\mu_{2A}(t)$. For small values of $t$, we expand both $a(t)$ and $\mu_{1A}(t)$ in a Taylor series around $t=0$,
@@
Hence, provided $N\geq\lfloor\tilde{M}/2\rfloor$, we have
@@
$\ddot{\mu}_{1A}(0)=-k^{2}_{2}N(\tilde{M}-N)(2N-\tilde{M}-1).$
@@
$\tilde{M}/2<a(t_{i})<\mu_{1A}(t_{i})<\tilde{M}$, (41)
is satisfied, then necessarily $g(a(t_{i}))<g(\mu_{1A}(t_{i}))$. But since condition (41) holds for $t\gtrsim 0$ (see Eq. (38)), and given that ${k}_{2}\kappa_{2A}(t)>0$, it follows that
$\dot{a}(t_{i})<\dot{\mu}_{1A}(t_{i}).$ (42)
Therefore once the inequality (39) arises at $t\gtrsim 0$ (see Eq. 38), it is preserved as long as condition (41) remains satisfied.
The conditions (34) and (42) imply that ${a}(t_{i})<{\mu}_{1A}(t_{i})$, which explains the appearance of the stochastic delay as long as condition (41) is satisfied. Once $g(a)>g(\mu_{1A})$, which can occur when (41) is no longer fulfilled, inequality (42) may no longer hold. At a certain time $t_{de}$ equality is reached in Eq. (42), and for $t>t_{de}$ the inequality reverses its sign.
Nevertheless, we can still conclude that ${a}(t)\leq{\mu}_{1A}(t)$ for all times. This follows from the long-time limit $t\to\infty$, where $a(t)$ and $\mu_{1A}(t)$, as well as their derivatives, vanish. Combined with condition (32), which ensures that both curves decrease monotonically, the two trajectories cannot intersect, confirming the inequality for all $t$.
The above analysis also reveals a close connection between the fluctuations, as quantified by the variance $\kappa_{2A}(t)=\sigma^{2}(t)$ or standard deviation $\sigma(t)$, and the difference between the time derivatives of $a(t)$ and $\mu_{1A}(t)$, and hence with the stochastic delay.
This connection is clearly visible in our numerical results presented so far (see Figs. 14 and Fig. 6), and also follows from the form of the r.h.s. of Eq. (29). Nevertheless, it can be further elucidated by analyzing Eq. (30).
@@
All these semiquantitative predictions are borne out by the numerical results for both the CME solution and the moment-equation solution. For $\kappa_{3A}(t)=0$ and $N_{B}(0)=1$, the solution of Eqs. (29)–(30) is shown in Fig. 7. Note that significant quantitative discrepancies are observed between this approximate solution (green) and the exact result (i.e., the solution of the CME, blue). For $N_{B}(0)=5$ (not shown), the approximate and exact stochastic solutions are essentially indistinguishable. However, in the most interesting case, $N_{B}(0)=0$, we were unable to obtain a stable numerical solution of Eqs. (29)–(30): despite employing various methods, including those designed for stiff systems, the solution invariably diverged.
@@ III.3 Two Reaction Networks Closely Related to the Finke–Watzky Model
From a mathematical perspective, the solution of the CME for each of the three models considered here can be obtained from that of the others by an appropriate mapping of model parameters. However, from the viewpoint of properties relevant to chemistry or population dynamics it is appropriate to treat the original Finke–Watzky model (1)–(2) and the inverse Finke–Watzky model (58)–(59) on an equal footing: they belong to distinct classes, one exhibiting pronounced stochastic delay and the other not. The reaction network (60)–(61), on the other hand, belongs, depending on the choice of model parameters, either to the same class as (1)–(2) or to that of (58)–(59). Therefore, it is convenient to view the solution for (60)–(61) as derived either from the solution of the FWM [Eqs. (1)–(2)] or from the solution of the inverse FWM [Eqs. (58)–(59)], depending on $\chi$ and $M$.
@@
$\tfrac{1}{2}\,k_{a}(\chi-1)=k_{2},\quad\frac{M\chi-1}{\chi-1}=\tilde{M}$. (71)
@@ IV Summary and Discussion
This work represents a generalization of the preliminary study of the stochastic delay effect Schuster (2019). Our findings further suggest that a pronounced stochastic delay—much larger than in the case of simple autocatalysis analyzed in Ref. Schuster (2019)—occurs not only in the Finke–Watzky model (1)–(2), but also in other reaction networks. In particular, substantial discrepancies between stochastic and deterministic kinetics are expected in networks that combine a first-order reaction, $\mathrm{A}\rightarrow\mathrm{B}$, with simple or higher-order autocatalysis (e.g., $2\mathrm{A}+\mathrm{B}\rightarrow\mathrm{A}+2\mathrm{B}$ or $\mathrm{A}+2\mathrm{B}\rightarrow 3\mathrm{B}$). [^7] For the reaction networks $\{\mathrm{A}\rightarrow\mathrm{B},2\mathrm{A}+\mathrm{B}\rightarrow\mathrm{A}+2\mathrm{B}\}$ and $\{\mathrm{A}\rightarrow\mathrm{B},\mathrm{A}+2\mathrm{B}\rightarrow 3\mathrm{B}\}$, this has indeed been confirmed by our preliminary numerical results.
@@ Acknowledgments
The authors would like to thank Aleksandra Siklitckaia, Adam Kubas, Sanat Kumar Mahankudo, and Gonzalo Angulo. TB was supported by the National Science Centre Grant Opus 20 Project No. 2020/39/B/ST4/01952.
@@ C.2 Reversible generalization of the Finke-Watkzy model (3)–(4)
"Deterministic" kinetic reaction rate constants $\mathcal{K}^{\pm}_{1}$, $\mathcal{K}^{\pm}_{2}$ appearing in the above equations correspond to the "stochastic" rate constants $k^{\pm}_{1}$ and $k^{\pm}_{2}$ of the CME (13) with $r_{n}$ (44) and $g_{n}$ (45), see Eq. (46). As in the case of Eq. (90)-(91), here we again have $a(t)+b(t)=a(0)+b(0)\equiv\mathcal{M}$. Therefore instead of (95)-(96) we obtain a single equation
@@ C.3 The inverse Finke–Watzky model (58)–(59)
Importantly, the relationship between deterministic and stochastic rate constants differs from that in the original FWM. In the present case we have

### v4

@@ I Introduction
For all of the above phenomena, the first-order process (1) is indispensable, and therefore autocatalysis (2) alone is not sufficient as a minimal effective model. [^2] An assessment of how realistic the Finke–Watzky model is for describing the phenomena and processes discussed here, and to what extent it can serve as their minimal, effective quantitative description, lies beyond the scope of the present work; see, however, Ref. Jędrak (2013); Jedrak (2014).
The deterministic rate equations for reactions (1)–(2) are straightforward to solve and have been used to fit experimental data on nanoparticle formation Watzky and Finke (1997) as well as on protein misfolding and aggregation in vitro Watzky et al. (2008); Morris et al. (2008, 2009); Iashchishyn et al. (2017). However, some applications of the FWM may involve relatively small systems—for example, modeling protein misfolding and aggregation in vivo or disease spread in small populations—where stochastic approaches such as the CME or SSA are more appropriate. The existence of significant differences between stochastic and deterministic descriptions of this model further motivates deriving an analytical solution of the CME.
@@ II.1 Irreversible reactions
The solution (10)–(11) of Eqs. (6) has appeared in various contexts and has been derived independently by different authors Allnatt and Jacobs (1968); Arslan and Laurenzi (2008); Lee and Kim (2012); Rudnicki (2014); Jedrak (2014). In contrast, by "explicit analytical solution" we mean here not the general form (10)–(11) of the solution, but the most concise analytical form of both $P_{n}(t)$ and its lowest moments that can be obtained for a specific choice of the coefficients $r_{n}$.
@@ III.1.1 Probability distribution and its two lowest moments
Analytical expressions for the moments (III.1.1)–(III.1.1) reduce the double sum (which arises when calculating the moment directly from its definition using $P_{n}(t)$ given by Eq. 22) to a single sum, lowering the computational cost. A caveat is that numerical evaluation of Eqs. (III.1.1) and (III.1.1) requires arbitrary-precision arithmetic: for $M=100$ the absolute values of the coefficients $E^{(1)}_{l}$ and $E^{(2)}_{l}$ reach $10^{30}$ and $10^{33}$, respectively, so small relative round-off errors at short times can cause large absolute errors in the moments. [^4] Here, the GNU Multiple Precision Floating-Point Reliable Library (GNU MPFR) was used for this purpose (https://www.mpfr.org). In fact, instead of relying on the potentially cumbersome arbitrary-precision evaluation of Eqs. (III.1.1)–(III.1.1), one can simply solve the set of $N+1$ ODEs in Eq. (6) by direct numerical integration. We performed the calculations using both approaches and obtained perfect agreement.
@@ III.1.3 Qualitative explanation of the observed effects
To study the kinetics of the Finke–Watzky model, one may replace the full CME analysis with an approximate approach based on a truncated system of moment equations. Here we examine such equations in order to provide a partial, semiquantitative explanation for the stochastic delay phenomenon. At the same time, our results highlight the limitations of moment-equation approaches when it comes to accurate quantitative prediction of the magnitude of differences between stochastic and deterministic kinetics.
@@
where $\tilde{M}=M+k_{1}/k_{2}$ (see Eq. (20)). Equations (29) and (31) imply that
@@
The former inequality is immediate (as $0\leq a\leq\tilde{M}$), while the latter follows from the relations $\tilde{M}\,\mu_{1A}(t)\geq M\,\mu_{1A}(t)\geq N\,\mu_{1A}(t)$ and $N\,\mu_{1A}(t)\geq\mu_{2A}(t)={\kappa}_{2A}(t)+\mu^{2}_{1A}(t)$.
For small values of $t$, we expand both $a(t)$ and $\mu_{1A}(t)$ in a Taylor series around $t=0$,
@@
Hence, provided $N\geq\lfloor\tilde{M}/2\rfloor$, by comparing the time derivatives of Eqs. (29) and (31), we obtain
@@
$\ddot{\mu}_{1A}(0)=-k^{2}_{2}N(\tilde{M}-N)(2N-\tilde{M}-1)<0.$
@@
$\tilde{M}/2<a(t)<\mu_{1A}(t)<\tilde{M}$, (41)
is satisfied, then necessarily $g(a(t))<g(\mu_{1A}(t))$. But since condition (41) holds for $t\gtrsim 0$ (see Eq. (38)), and given that ${k}_{2}\kappa_{2A}(t)>0$, from Eqs. (29) and (31) it follows that
$\dot{a}(t)<\dot{\mu}_{1A}(t).$ (42)
Therefore once the inequality (39) arises at $t\gtrsim 0$ (see Eq. 38), it is preserved as long as condition (41) remains satisfied. The conditions (34) and (42) imply that ${a}(t)<{\mu}_{1A}(t)$, which explains the appearance of the stochastic delay as long as condition (41) is satisfied. But once $g(a)>g(\mu_{1A})$, which can happen when (41) is not satisfied, inequality (42) may no longer hold. At a certain time $t=t_{de}$ equality is reached in Eq. (42), and for $t>t_{de}$ the inequality reverses its sign.
Nevertheless, we can still conclude that ${a}(t)<{\mu}_{1A}(t)$ for all times. Suppose that, for some $t=t_{fe}$, at which condition (41) is no longer satisfied, the curves a(t){a}(t) and ${\mu}_{1A}(t)$ intersect for the first time, i.e., ${a}(t_{fe})={\mu}_{1A}(t_{fe})$. Since ${a}(t)\leq{\mu}_{1A}(t)$ for all $t<t_{fe}$, this would require $\dot{\mu}_{1A}(t_{fe})<\dot{a}(t_{fe})$. However, this is impossible: a comparison of Eqs. (29) and (31) shows that whenever ${a}(t_{fe})={\mu}_{1A}(t_{fe})$, one must have $\dot{\mu}_{1A}(t_{fe})\geq\dot{a}(t_{fe})$, with equality only if $\kappa_{2A}(t_{fe})=0$. This confirms that ${a}(t)\leq{\mu}_{1A}(t)$ (and thus the stochastic delay) for all $t$.
The above analysis also reveals a close connection between the fluctuations, as quantified by the variance $\kappa_{2A}(t)=\sigma^{2}(t)$ or standard deviation $\sigma(t)$, and the difference between the time derivatives of $a(t)$ and $\mu_{1A}(t)$, and hence with the stochastic delay. This connection is clearly visible in our numerical results presented so far (see Figs. 14 and Fig. 6), and also follows from the form of the r.h.s. of Eq. (29). Nevertheless, it can be further elucidated by analyzing Eq. (30).
@@
All these semiquantitative predictions are borne out by the numerical results for both the CME solution and the moment-equation solution. For $\kappa_{3A}(t)=0$ and $N_{B}(0)=1$, the solution of Eqs. (29)–(30) is shown in Fig. 7. Note that significant quantitative discrepancies are observed between this approximate solution (green) and the exact result (i.e., the solution of the CME, blue). For $N_{B}(0)=5$ (not shown), the approximate and exact stochastic solutions are essentially indistinguishable. However, in the most interesting case, $N_{B}(0)=0$, we were unable to obtain a stable numerical solution of Eqs. (29)–(30): despite using various methods, including those for stiff systems, and different choices of dependent variables, the solution invariably diverged.
@@ III.3 Two Reaction Networks Closely Related to the Finke–Watzky Model
From a mathematical perspective, the solution of the CME for each of the three models considered here can be obtained from that of the others by an appropriate mapping of model parameters. However, from the viewpoint of properties relevant to chemistry or population dynamics it is appropriate to treat the original Finke–Watzky model (1)–(2) and the inverse Finke–Watzky model (58)–(59) on an equal footing: they belong to distinct classes, one exhibiting pronounced stochastic delay and the other not. The reaction network (60)–(61), on the other hand, belongs, depending on the choice of model parameters, either to the same class as (1)–(2) or to that of (58)–(59).
Therefore, it is convenient to view the solution for (60)–(61) as derived either from the solution of the FWM [Eqs. (1)–(2)] or from the solution of the inverse FWM [Eqs. (58)–(59)], depending on $\chi$ and $M$.
@@
$\tfrac{1}{2}\,k_{a}(\chi-1)=k_{2},\quad\frac{M\chi-1}{\chi-1}=\tilde{M}>M.$ (71)
Thus, for large values of $\chi$ ($k_{a}\ll k_{b}$), we have $\tilde{M}\gtrsim M$, and by analogy with the original Finke–Watzky model, a significant stochastic delay can be expected when the system is initialized with no B molecules. This is evident even without detailed analysis: in this parameter regime, reaction (60) effectively plays the role of the first-order process (1) in the original Finke–Watzky model, providing a low-probability “nucleation” or “seeding” event that is subsequently amplified by autocatalysis (61).
@@
It is evident that in the limit $k_{a}\gg k_{b}$ , the behavior of the system described by Eqs. (60)–(61) is predominantly governed by the first of these two reactions.
@@ IV Summary and Discussion
This work represents a generalization of the preliminary study of the stochastic delay effect Schuster (2019). Our findings further suggest that a pronounced stochastic delay—much larger than in the case of simple autocatalysis analyzed in Ref. Schuster (2019)—occurs not only in the Finke–Watzky model (1)–(2), but also in other reaction networks. In particular, substantial discrepancies between stochastic and deterministic kinetics are expected in networks that combine a first-order reaction, $\mathrm{A}\rightarrow\mathrm{B}$, with higher-order autocatalysis (e.g., $2\mathrm{A}+\mathrm{B}\rightarrow\mathrm{A}+2\mathrm{B}$ or $\mathrm{A}+2\mathrm{B}\rightarrow 3\mathrm{B}$). [^8] For the reaction networks $\{\mathrm{A}\rightarrow\mathrm{B},2\mathrm{A}+\mathrm{B}\rightarrow\mathrm{A}+2\mathrm{B}\}$ and $\{\mathrm{A}\rightarrow\mathrm{B},\mathrm{A}+2\mathrm{B}\rightarrow 3\mathrm{B}\}$, this has indeed been confirmed by our preliminary numerical results.
@@
Data Availability
All raw data and source codes used to generate the figures in this manuscript are openly available in Ref. Jędrak (2025).
@@ Acknowledgments
The authors would like to thank Aleksandra Siklitckaia, Adam Kubas, Sanat Kumar Mahankudo, Marcin Rubin, and Gonzalo Angulo. TB was supported by the National Science Centre Grant Opus 20 Project No. 2020/39/B/ST4/01952.
@@ C.2 Reversible generalization of the Finke-Watkzy model (3)–(4)
"Deterministic" kinetic reaction rate constants $\mathcal{K}^{\pm}_{1}$, $\mathcal{K}^{\pm}_{2}$ appearing in the above equations correspond to the "stochastic" rate constants $k^{\pm}_{1}$ and $k^{\pm}_{2}$ of the CME (13) with $r_{n}$ (44) and $g_{n}$ (45), see Eq. (46). As in the case of Eq. (90)-(91), here we again have $a(t)+b(t)=a(0)+b(0)\equiv\mathcal{M}$. Therefore instead of (95)-(96) we can use a single equation
@@ C.3 The inverse Finke–Watzky model (58)–(59)
The relationship between deterministic and stochastic rate constants differs from that in the original FWM. In the present case we have